# Microphysical Regulation of Non-Ideal MHD in Weakly-Ionized Systems: Does the Hall Effect Matter?

### v1

@@ 3 Where Might the “Standard” Formulation Be Problematic?
Where might the correction terms in § 2.4 be important? Consider the case with $v_{\rm drift}>v_{\rm slip}$, and note we can write $\ell^{\rm drift}_{\rm crit}\approx(v_{A,i}/v_{T})d_{i}$, where $v_{A,i}\equiv B/\sqrt{4\pi\rho_{i}}$ and $d_{i}\equiv c/\omega_{pi}=v_{A,i}/\Omega_{c,i}=r_{g,i}/\sqrt{\beta^{\rm plasma}_{i}}$ are the ion Alfvén speed and inertial length (and $\beta^{\rm plasma}_{i}\equiv(c_{s}/v_{A,\,i})^{2}\equiv P_{{\rm therm},\,i}/2\,P_{B}\equiv 4\pi\,n_{i}\,k_{B}\,T_{i}/B^{2}\equiv f_{i}\,\beta^{\rm plasma}$ is the ion plasma $\beta^{\rm plasma}_{i}$, with $r_{g,i}$ the ion gyro radius). This means that for well-ionized plasmas, so long as $\ell_{B}\gg d_{i}$ (for $v_{T}\sim v_{A}$) or $\gg d_{i}/\beta_{\rm plasma}^{1/2}=r_{g}/\beta_{\rm plasma}$ (for $v_{T}\sim v_{\rm thermal}$), the system will be “safely” in the $v_{\rm drift}\ll v_{T,\,i}$ regime. [^5] For example, in the diffuse interstellar, circumgalactic, and intergalactic medium, stellar photospheres, HII regions, and OB winds, the ratio of typical $\ell_{B}$ for order-unity changes in $B$ (i.e. the Alfvén scale of turbulence) to $\ell_{\rm crit}$ is $\sim 10^{10}-10^{15}$, and it reaches $\gg 10^{20}$ in stellar interiors. Even in extremely low-$\beta^{\rm plasma}$ environments (as low as $<10^{-6}$), such as those proposed for quasar accretion disks in Hopkins et al. (2023a, c, b) this ratio is still $\gg 10^{9}$.
@@ 4.1 Ordering and Scaling of Terms
The regime where ambipolar diffusion would dominate, in the “standard” or zero-drift-assumption limit ($\Omega_{c,\pm}\gg\omega^{0}_{\rm coll}$), generically has $v_{\rm drift}\ll v_{\rm slip}$, meaning current-driven instabilities will not be important. While there can be pure slip-driven instabilities, their effect is less clear (see § 2.3.2). So when the slip becomes superthermal, we just have the Epstein-type correction for $\langle\sigma v\rangle$ accounting for the superthermal relative velocity of charge carriers versus neutrals in the collision rates, which decreases $\eta_{A}$ and increases $\eta_{O}$, but only by a linear factor of $v_{\rm drift/slip}/v_{T}$. Thus it will not usually change the relative ordering of the non-ideal terms.
In the regime where Ohmic resistivity would strongly dominate with zero drift ($\Omega_{c,\pm}\ll\omega_{\rm coll}$), $v_{\rm drift}\gg v_{\rm slip}$. If the drift becomes superthermal, current-driven instabilities could appear but in this limit, the anomalous resistivity $\eta_{O}^{\rm an}\sim\eta_{O}^{0}(\Omega_{c-,p-}/\omega_{\rm coll}^{0})\ll\eta_{O}^{0}$ would be relatively small because the collision frequency driving the resistivity is already the fastest frequency of interest (so even if collisions strongly damp these instabilities in this limit, it makes no practical difference). The dominant correction will again simply be the proper accounting for superthermal drift in the Epstein correction to the collision rates, boosting $\eta_{O}$ by a factor of $\sim v_{\rm drift}/v_{T}$.
The regime where the Hall term would dominate with zero drift ($\Omega_{c,+}\ll\omega_{\rm coll}\ll\Omega_{c,-}$, $v_{\rm drift}\sim v_{\rm slip,-}\gg v_{\rm slip,+}$) is perhaps most interesting, despite the fact that the Hall term itself is independent of any of these correction factors (being independent of the effective collision frequency except for an $\mathcal{O}(1)$ prefactor that depends only on the dimensionless ratios of different collision rates). In this regime, by definition, the zero-drift Ohmic term is relatively small $\eta_{O}^{0}\sim|\eta_{H}^{0}|\,\omega_{\rm coll}/\Omega_{c-}\ll|\eta_{H}^{0}|$. But if the drift becomes superthermal, current-driven instabilities will appear and be weakly or negligibly damped. Thus $\eta_{O}\rightarrow(\eta_{O}^{0}+\eta_{O}^{\rm an})\,(v_{\rm drift}/v_{T})$, where $\eta_{O}^{\rm an}\sim|\eta_{H}|$ and $v_{\rm drift}/v_{T}\gtrsim 1$. So $\eta_{O}\sim|\eta_{H}|\,(v_{\rm drift}/v_{T})\gtrsim|\eta_{H}|$ and the excited Ohmic term always becomes larger than the Hall term. The outcome is that one can never have a dominant Hall term and superthermal drift speeds (see also Choueiri, 1999; Shay et al., 1999).
@@ 4.2 Driving the System Towards Sub-Thermal Drift Speeds and Smooth Magnetic Gradients
Consider what happens if superthermal drift arises, i.e. something induces gradients with $\ell_{B}\ll\ell_{\rm crit}$. As discussed in § 3, if we only adopt the zero-drift coefficients $\eta_{O,H,A}^{0}$, then especially if the Hall term dominates, there is no guarantee that the non-ideal terms will actually reduce the drift speed (or increase $\ell_{B}$) because it is non-diffusive. But per § 4.1, we see that with the corrected coefficients, $\eta_{H}$ can never dominate in this limit, so we are ensured that the nonideal terms behave diffusively (see Fig. 2). This means that they will always act to smooth out the gradients $\ell_{B}$ until subthermal drift is restored. The timescale for this to occur will be $\sim\ell_{B}^{2}/{\rm MAX}[\eta_{O},\,\eta_{A}]\sim(\ell_{B}/v_{T})\,(v_{\rm drift}/v_{T})^{-2}\,\mathcal{F}_{\rm max}^{-1}$ where $v_{\rm drift}/v_{T}\gtrsim 1$ and $\mathcal{F}_{\rm max}\equiv{\rm MAX}[1,\,\Omega_{+}/\omega_{\rm coll},\,\omega_{p-}/\Omega_{-},\,\omega_{\rm coll}/\Omega_{-}]\gtrsim 1$. So generically this will occur on a timescale much shorter than the sound crossing time.
@@
Note here that $\ell_{B}\equiv|\bar{\bf B}|/|\nabla\times\bar{\bf B}|$ is the global gradient length scale; so weak magnetic field gradients on scales $\lambda\ll\ell_{\rm crit}$ are still allowed. If we consider some small-amplitude structure with |×$|\nabla\times\bar{\bf B}|\sim\delta B/\lambda$ and $\delta B\ll B$; then $\ell_{B}\sim\lambda\,|B|/|\delta B|\gtrsim\ell_{\rm crit}$. In this limit we can thus think of the physics above as setting an upper limit to the strength of field perturbations on scale $\lambda$, i.e. $|\delta B|/|B|\lesssim\lambda/\ell_{\rm crit}$. This implies that the modification to MHD waves from the Hall effect will remain, but that these waves will experience non-linear damping above a certain (wavelength-dependent) amplitude owing to induced super-thermal drift enhancing the Ohmic resistivity. Recalling that $\ell_{\rm crit}^{\rm drift}\sim r_{g}/(f_{\rm ion}\beta^{\rm plasma})$, this becomes $|\delta B|/|B|\lesssim(\lambda/r_{g})\,f_{\rm ion}\,\beta^{\rm plasma}$, we see that this places a serious constraint in very low $f_{\rm ion}$ plasmas for short-wavelength MHD waves and other perturbations. This is similar to effects known in high-$\beta^{\rm plasma}$ collisionless plasmas due to analogous instability-mediated “anomalous viscosities”: Alfvén waves cannot propagate (Squire et al., 2017c) and sound waves are modified (Kunz et al., 2014) above a $\beta^{\rm plasma}$-dependent threshold.
@@ 4.5 Is Non-Ideal Hall MHD Ever Important?
From the above (§ 4.1 & Figs. 1-2), the non-ideal Hall term can only dominate in the induction equation when $\Omega_{+}\ll\omega_{\rm coll}\ll\Omega_{-}$ and $|{\bf U}|\ll v_{\rm drift}\ll v_{T}$. This is not theoretically impossible; it is however challenging in practice to imagine in equilibrium on large scales. For example, it is only possible if $|{\bf U}|\ll v_{T,\,i}$, i.e. if we have some flow with negligible gas (sub-thermal) gas motions on some relevant global scale $\ell_{B}$, within some modest window of density and field strength (Fig. 1), with $\ell_{B}\gg\ell_{\rm crit}$.
[…]
@@
For weak perturbations, as noted above (§ 4.2), the Hall term could still be important. Consider an incompressible magnetic fluctuation which is also weak (amplitude $\delta B\ll(\lambda/\ell_{\rm crit})\,|B|$) and on sufficiently short wavelengths $\lambda\ll\ell_{\rm crit}\ll\ell_{B}$ such that we can boost to a frame moving with the mean ${\bf U}$ and have only small fluctuations $\delta{\bf U}(\lambda)$ on scale $\lambda$. If we are in the limit where the “classical” Hall term is larger than the Ohmic and ambipolar terms, then in this limit we will just recover the usual linearized Hall-MHD equations: i.e. that the Hall term modifies short-wavelength Alfvén waves to ion-cyclotron and whistler waves. So our intuition for these is unchanged, except that when we consider finite-amplitude waves, their amplitude must remain very small with $|\delta B|\ll(\lambda/\ell_{\rm crit})\,|B|$ to avoid exciting current-driven instabilities.
@@ 4.6 Behavior in the Well-Ionized Limit
In this limit, then, the Hall term can become important on some (small) scale $\lambda$ so long as $|\delta{\bf U}(\lambda)|\lesssim v_{\rm drift}\lesssim v_{T}$. In the fully ionized limit the drift velocities are small, $v_{\rm drift}\sim v_{T}\ell_{\rm crit}/\ell_{B}\sim(v_{T}/\beta^{\rm plasma})\,(r_{g}/\ell_{B})\sim(v_{T}/\beta^{\rm plasma})\,(r_{g}/\lambda)\,(|\delta B|/|B|)$. Thus the Hall term being important requires $|\delta{\bf U}(\lambda)|/v_{T}\ll(r_{g}/\beta^{\rm plasma}\lambda)\,(|\delta B|/|B|)\ll 1$. Since MHD is only valid on scales $\lambda\gg r_{g}$, the latter inequality ($v_{\rm drift}\ll v_{T}$ being sub-thermal) is essentially always satisfied, while the former inequality ($|\delta{\bf U}(\lambda)|\ll v_{\rm drift}$) is valid on sufficiently small scales $\lambda$. Note that on large scales in ISM and other applications, $|\delta{\bf U}(\lambda)|/v_{T}\gtrsim 1$, i.e. bulk motions are trans or super-sonic, and $\lambda\gg r_{g}$ by a huge factor while $|\delta B|\lesssim|B|$, which means that the Hall term is completely negligible compared to the ideal MHD term ${\bf U}\times{\bf B}$ – this is just a restatement of the fact that ideal MHD is an excellent approximation on large scales in well-ionized astrophysical plasmas. But this again leads to the usual expected small-scale behaviors that Hall terms will modify short-wavelength Alfvén waves as expected in a well-ionized plasma, though other effects (e.g. electron inertia) can also become important on small scales (Schekochihin et al., 2009).
@@ 5 Conclusions
The commonly-adopted astrophysical MHD equations are no longer internally self-consistent if gradients in ${\bf B}$ become too steep (gradient scale-length $\ell_{B}$ smaller than some critical $\ell_{\rm crit}$), because the implied drift speeds ($v_{\rm drift}$) become much faster than thermal speeds ($v_{T}$). This can invalidate foundational assumptions of the equations, rendering the expressions generally adopted for certain terms (like the non-ideal Ohmic/Hall/ambipolar coefficients) incorrect. Most importantly, it will excite microscale instabilities that suppress the drift, potentially modifying the effective collision frequencies by orders of magnitude. We show that this could be relevant in weakly ionized systems like cold cores, circumstellar/planetary disks, or planetary atmospheres, especially if (1) the Hall effect is significant, (2) external forces (like gravity) induce highly supersonic motions, or (3) dust grains become an important charge carrier. The “classical” treatments could then predict unphysical behaviors. We derive a simple prescription to remedy this, extending treatments well-established in the fusion, Solar, and ionospheric plasma literature, and directly analogous to well-established treatments for similar problems in thermal conduction and viscosity. The treatment amounts to simply multiplying the Ohmic and ambipolar coefficients by factors involving only macroscopic (already-calculated) quantities. Implementing these correction terms in simulations therefore entails negligible cost or complexity. They also leave the behavior unchanged in the limit of slow drift for which the classical expressions are derived, but will restore the physical behavior if this limit is violated, driving the system back towards slow drift.

### v2

@@ 3 Where Might the “Standard” Formulation Be Problematic?
Where might the correction terms in § 2.4 be important? Consider the case with $v_{\rm drift}>v_{\rm slip}$, and note we can write $\ell^{\rm drift}_{\rm crit}\approx(v_{A,i}/v_{T})d_{i}$, where $v_{A,i}\equiv B/\sqrt{4\pi\rho_{i}}$ and $d_{i}\equiv c/\omega_{pi}=v_{A,i}/\Omega_{c,i}=r_{g,i}/\sqrt{\beta^{\rm plasma}_{i}}$ are the ion Alfvén speed and inertial length (and $\beta^{\rm plasma}_{i}\equiv(c_{s}/v_{A,\,i})^{2}\equiv P_{{\rm therm},\,i}/2\,P_{B}\equiv 4\pi\,n_{i}\,k_{B}\,T_{i}/B^{2}\equiv f_{i}\,\beta^{\rm plasma}$ is the ion plasma $\beta^{\rm plasma}_{i}$, with $r_{g,i}$ the ion gyro radius). This means that for well-ionized plasmas, so long as $\ell_{B}\gg d_{i}$ (for $v_{T}\sim v_{A}$) or $\gg d_{i}/\beta_{\rm plasma}^{1/2}=r_{g}/\beta_{\rm plasma}$ (for $v_{T}\sim v_{\rm thermal}$), the system will be “safely” in the $v_{\rm drift}\ll v_{T,\,i}$ regime. [^5] For example, in the diffuse interstellar, circumgalactic, and intergalactic medium, stellar photospheres, HII regions, and OB winds, the ratio of typical $\ell_{B}$ for order-unity changes in $B$ (i.e. the Alfvén scale of turbulence) to $\ell_{\rm crit}$ is $\sim 10^{10}-10^{15}$, and it reaches $\gg 10^{20}$ in stellar interiors. Even in extremely low-$\beta^{\rm plasma}$ environments (as low as $10^{-6}$), such as those proposed for quasar accretion disks in Hopkins et al. (2023a, c, b) this ratio is still $\gg 10^{9}$.
@@ 4.1 Ordering and Scaling of Terms
The regime where ambipolar diffusion would dominate, in the “standard” or zero-drift-assumption limit ($\Omega_{c,\pm}\gg\omega^{0}_{\rm coll}$), generically has $v_{\rm drift}\ll v_{\rm slip}$, meaning current-driven instabilities will not be important. While there can be pure slip-driven instabilities, their effect is less clear (see § 2.3.2). So when the slip becomes superthermal, we just have the Epstein-type correction for $\langle\sigma v\rangle$ accounting for the superthermal relative velocity of charge carriers versus neutrals in the collision rates, which decreases $\eta_{A}$ and increases $\eta_{O}$, but only by a linear factor of $v_{\rm drift/slip}/v_{T}$. Thus it will not usually change the relative ordering of the non-ideal terms (unless they are already similar in magnitude).
In the regime where Ohmic resistivity would strongly dominate with zero drift ($\Omega_{c,\pm}\ll\omega_{\rm coll}$), $v_{\rm drift}\gg v_{\rm slip}$. If the drift becomes superthermal, current-driven instabilities could appear but in this limit, the anomalous resistivity $\eta_{O}^{\rm an}\sim\eta_{O}^{0}(\Omega_{c-,p-}/\omega_{\rm coll}^{0})\ll\eta_{O}^{0}$ would be relatively small because the collision frequency driving the resistivity is already the fastest frequency of interest. So even if collisions strongly damp these instabilities in this limit, it makes no practical difference. The dominant correction will again simply be the proper accounting for superthermal drift in the Epstein correction to the collision rates, boosting $\eta_{O}$ by a factor of $\sim v_{\rm drift}/v_{T}$.
The regime where the Hall term would dominate with zero drift ($\Omega_{c,+}\ll\omega_{\rm coll}\ll\Omega_{c,-}$, $v_{\rm drift}\sim v_{\rm slip,-}\gg v_{\rm slip,+}$) is perhaps most interesting, despite the fact that the Hall term itself is independent of these correction factors to leading order (being independent of the effective collision frequency except for an $\mathcal{O}(1)$ prefactor that depends only on the dimensionless ratios of different collision rates). In this regime, by definition, the zero-drift Ohmic term is relatively small $\eta_{O}^{0}\sim|\eta_{H}^{0}|\,\omega_{\rm coll}/\Omega_{c-}\ll|\eta_{H}^{0}|$. But if the drift becomes superthermal, current-driven instabilities will appear and be weakly or negligibly damped. Thus $\eta_{O}\rightarrow(\eta_{O}^{0}+\eta_{O}^{\rm an})\,(v_{\rm drift}/v_{T})$, where $\eta_{O}^{\rm an}\sim|\eta_{H}|$ and $v_{\rm drift}/v_{T}\gtrsim 1$. So $\eta_{O}\sim|\eta_{H}|\,(v_{\rm drift}/v_{T})\gtrsim|\eta_{H}|$ and the excited Ohmic term generically becomes larger than the Hall term. The outcome is that one cannot have a dominant Hall term and superthermal drift speeds (see also Choueiri, 1999; Shay et al., 1999).
@@ 4.2 Driving the System Towards Sub-Thermal Drift Speeds and Smooth Magnetic Gradients
Consider what happens if superthermal drift arises, i.e. something induces gradients with $\ell_{B}\ll\ell_{\rm crit}$. As discussed in § 3, if we only adopt the zero-drift coefficients $\eta_{O,H,A}^{0}$, then especially if the Hall term dominates, there is no guarantee that the non-ideal terms will actually reduce the drift speed (or increase $\ell_{B}$) because it is non-diffusive. But per § 4.1, we see that with the corrected coefficients, $\eta_{H}$ does not dominate in this limit, so we are ensured that the nonideal terms behave diffusively (see Fig. 2). This means that they will always act to smooth out the gradients $\ell_{B}$ until subthermal drift is restored. The timescale for this to occur will be $\sim\ell_{B}^{2}/{\rm MAX}[\eta_{O},\,\eta_{A}]\sim(\ell_{B}/v_{T})\,(v_{\rm drift}/v_{T})^{-2}\,\mathcal{F}_{\rm max}^{-1}$ where $v_{\rm drift}/v_{T}\gtrsim 1$ and $\mathcal{F}_{\rm max}\equiv{\rm MAX}[1,\,\Omega_{+}/\omega_{\rm coll},\,\omega_{p-}/\Omega_{-},\,\omega_{\rm coll}/\Omega_{-}]\gtrsim 1$. So generically this will occur on a timescale much shorter than the sound crossing time.
@@
Note here that $\ell_{B}\equiv|\bar{\bf B}|/|\nabla\times\bar{\bf B}|$ is the global gradient length scale; so weak magnetic field gradients on scales $\lambda\ll\ell_{\rm crit}$ are still allowed. If we consider some small-amplitude structure with $|\nabla\times\bar{\bf B}|\sim\delta B/\lambda$ and $\delta B\ll B$; then sub-thermal drift requires $\ell_{B}\sim\lambda\,|B|/|\delta B|\gtrsim\ell_{\rm crit}$. In this limit we can thus think of the physics above as setting an upper limit to the strength of field perturbations on scale $\lambda$, i.e. $|\delta B|/|B|\lesssim\lambda/\ell_{\rm crit}$. This implies that the modification to MHD waves from the Hall effect will remain, but that these waves will experience non-linear damping above a certain (wavelength-dependent) amplitude owing to induced super-thermal drift enhancing the Ohmic resistivity. Recalling that $\ell_{\rm crit}^{\rm drift}\sim r_{g}/(f_{\rm ion}\beta^{\rm plasma})$, this becomes $|\delta B|/|B|\lesssim(\lambda/r_{g})\,f_{\rm ion}\,\beta^{\rm plasma}$. We see that this places a serious constraint in very low $f_{\rm ion}$ plasmas for short-wavelength MHD waves and other perturbations. This is similar to effects known in high-$\beta^{\rm plasma}$ collisionless plasmas due to analogous instability-mediated “anomalous viscosities”: Alfvén waves cannot propagate (Squire et al., 2017c) and sound waves are modified (Kunz et al., 2014) above a $\beta^{\rm plasma}$-dependent threshold.
@@ 4.5 Is Non-Ideal Hall MHD Ever Important?
From the above (§ 4.1 & Figs. 1-2), the non-ideal Hall term can only strongly dominate in the induction equation when $\Omega_{+}\ll\omega_{\rm coll}\ll\Omega_{-}$ and $|{\bf U}|\ll v_{\rm drift}\ll v_{T}$. This is not theoretically impossible; it is however challenging in practice to imagine in equilibrium on large scales. For example, it is only possible if $|{\bf U}|\ll v_{T,\,i}$, i.e. if we have some flow with negligible gas (sub-thermal) gas motions on some relevant global scale $\ell_{B}$, within some modest window of density and field strength (Fig. 1), with $\ell_{B}\gg\ell_{\rm crit}$.
Note that the Hall factor $(\nabla\times{\bf B})\times{\bf B}$ is the same term that appears in the total momentum equation: so for the Hall term to be important in induction on large scales (e.g. to drive formation of current sheets or magnetic switches, or “torques up” the disk, as observed in the Hall-dominated regime using the zero-drift coefficients), it must necessarily appear in the momentum equation as a magnetic pressure or tension force. Further it is easy to verify that the interesting Hall terms generically correspond to compressible fluctuations ($\nabla\cdot[(\nabla\times{\bf B})\times{\bf B}]\neq 0$). So we should compare both restoring magnetic-tension forces (with characteristic timescales $\sim\ell_{B}/v_{A}$) and pressure forces (with timescales $\sim\ell_{B}/v_{T}$), to the characteristic Hall induction timescale $\sim\ell_{B}^{2}/|\eta_{H}|$. Noting the scaling of $|\eta_{H}|\sim v_{\rm drift}\ell_{B}$, the Hall timescale is $\sim\ell_{B}/v_{\rm drift}$, so can only be shorter than thermal timescales if $v_{\rm drift}\gg v_{T}$, but in that superthermal limit the Ohmic term always becomes larger than Hall. So only rather contrived Hall-driven effects would be expected to persist.
@@
For weak perturbations, as noted above (§ 4.2), the Hall term could still be important. Consider an incompressible magnetic fluctuation which is also weak (amplitude $|\delta B|\ll(\lambda/\ell_{\rm crit})\,|B|$) and on sufficiently short wavelengths $\lambda\ll\ell_{\rm crit}\ll\ell_{B}$ such that we can boost to a frame moving with the mean $\bf U$} and have only small fluctuations $\delta{\bf U}(\lambda)$ on scale $\lambda$. If we are in the limit where the “classical” Hall term is larger than the Ohmic and ambipolar terms, then in this limit we will just recover the usual linearized Hall-MHD equations: i.e. that the Hall term modifies short-wavelength Alfvén waves to ion-cyclotron and whistler waves. So our intuition for these is unchanged, except that when we consider finite-amplitude waves, their amplitude must remain very small with $|\delta B|\ll(\lambda/\ell_{\rm crit})\,|B|$ to avoid exciting current-driven instabilities.
@@ 4.6 Behavior in the Well-Ionized Limit
In this limit, then, the Hall term can become important on some (small) scale $\lambda$ so long as $|\delta{\bf U}(\lambda)|\lesssim v_{\rm drift}\lesssim v_{T}$. In the fully ionized limit the drift velocities are generally small, $v_{\rm drift}\sim v_{T}\ell_{\rm crit}/\ell_{B}\sim(v_{T}/\beta^{\rm plasma})\,(r_{g}/\ell_{B})\sim(v_{T}/\beta^{\rm plasma})\,(r_{g}/\lambda)\,(|\delta B|/|B|)$. Thus the Hall term being important requires $|\delta{\bf U}(\lambda)|/v_{T}\ll(r_{g}/\beta^{\rm plasma}\lambda)\,(|\delta B|/|B|)\ll 1$. Since MHD is only valid on scales $\lambda\gg r_{g}$, the latter inequality ($v_{\rm drift}\ll v_{T}$ being sub-thermal) is essentially always satisfied, while the former inequality ($|\delta{\bf U}(\lambda)|\ll v_{\rm drift}$) is valid on sufficiently small scales $\lambda$. Note that on large scales in ISM and other applications, $|\delta{\bf U}(\lambda)|/v_{T}\gtrsim 1$, i.e. bulk motions are trans or super-sonic, and $\lambda\gg r_{g}$ by a huge factor while $|\delta B|\lesssim|B|$, which means that the Hall term is completely negligible compared to the ideal MHD term ${\bf U}\times{\bf B}$. This is just a restatement of the fact that ideal MHD is an excellent approximation on large scales in well-ionized astrophysical plasmas. But this again leads to the usual expected small-scale behaviors that Hall terms will modify short-wavelength Alfvén waves as expected in a well-ionized plasma, though other effects (e.g. electron inertia) can also become important on small scales (Schekochihin et al., 2009).
@@ 5 Conclusions
The commonly-adopted astrophysical MHD equations are no longer internally self-consistent if gradients in $\bf B$} become too steep (gradient scale-length $\ell_{B}$ smaller than some critical $\ell_{\rm crit}$), because the implied drift speeds ($v_{\rm drift}$) become much faster than thermal speeds ($v_{T}$). This can invalidate foundational assumptions of the equations, rendering the expressions generally adopted for certain terms (like the non-ideal Ohmic/Hall/ambipolar coefficients) incorrect. Most importantly, it will excite microscale instabilities that suppress the drift, potentially modifying the effective collision frequencies by orders of magnitude. We show that this could be relevant in weakly ionized systems like cold cores, circumstellar/planetary disks, or planetary atmospheres, especially if (1) the Hall effect is significant, (2) external forces (like gravity) induce highly supersonic motions, or (3) dust grains become an important charge carrier. The “classical” treatments could then predict unphysical behaviors. We derive a simple prescription to remedy this, extending treatments established in the fusion, Solar, and ionospheric plasma literature, and directly analogous to well-established treatments for similar problems in thermal conduction and viscosity. The treatment amounts to simply multiplying the Ohmic and ambipolar coefficients by factors involving only macroscopic (already-calculated) quantities. Implementing these correction terms in simulations therefore entails negligible cost or complexity. They also leave the behavior unchanged in the limit of slow drift for which the classical expressions are derived, but will restore the physical behavior if this limit is violated, driving the system back towards slow drift.